HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SNAPS & HUBS

\underline{R. TÓBIÁS}, MTA-ELTE Complex Chemical Systems Research Group, H-1518 Budapest 112, P.O. Box 32, Hungary; A. G. CSÁSZÁR, Institute of Chemistry, ELTE Eötvös Loránd University, H-1117 Budapest, Pázmány Péter sétány 1/A, Hungary & MTA-ELTE Complex Chemical Systems Research Group, H-1518 Budapest 112, P.O. Box 32, Hungary; M. L. DIOUF, P. BLOEM, F. M. J.

COZIJN, W. UBACHS, Department of Physics and Astronomy, LaserLaB, Vrije Universiteit, De Boelelaan 1081, 1081 HV Amsterdam, The Netherlands

The theory of spectroscopic networks^{1} offers a powerful instrument for the intelligent design and validation of precision-spectroscopy experiments, as well as the subsequent derivation of accurate rovibrational energies,^{2}^{,}^{3}^{,}^{4}^{,}^{5} in particular those associated with hubs (high-degree nodes) of the network. Through a joint experimental and theoretical approach, absolute energies have been determined for a large number of hubs in the experimental spectroscopic networks of H_{2}^{16}O and H_{2}^{18}O, with an accuracy of a few times 10^{−7} cm^{−1}. The Lamb-dip measurements utilized two noise-immune cavity-enhanced optical heterodyne molecular spectroscopy (NICE-OHMS) apparata^{6}^{,}^{7} and three probe lasers covering different wavenumber ranges. The observed lines, whose lower states belong to the vibrational ground state and the bending fundamental, were selected via the spectroscopic-network-assisted precision spectroscopy (SNAPS) approach.2 The accurately known energy levels are involved in thousands of unique transitions already measured, though at much lower accuracy, for H_{2}^{16}O and H_{2}^{18}O. From the ultraprecise absolute energies a large number of benchmark-quality predicted transitions has been deduced, which could be employed as frequency standards in high-resolution Fourier-transform infrared spectroscopy.

1A. G. Császár and T. Furtenbacher. J. Mol. Spectrosc., 266, 99–103, (2011), DOI.

2 R. Tóbiás, T. Furtenbacher, I. Simkó, A. G. Császár, M. L. Diouf, F. M. J. Cozijn, J. M. A. Staa, E. J. Salumbides, and W. Ubachs, Nat. Commun., 11, 1708 (2020), DOI.

3M. L. Diouf, R. Tóbiás, I. Simkó, F. M. J. Cozijn, E. J. Salumbides, W. Ubachs, and A. G. Császár.

J. Phys. Chem. Ref. Data, 50, 023106 (2021) DOI.

4M. L. Diouf, R. Tóbiás, T. S. van der Schaaf, F. M. J. Cozijn, E. J. Salumbides, A. G. Császár, and W. Ubachs. Mol. Phys., 120, e2050430 (2022) DOI.

5R. Tóbiás, M. L. Diouf, P. Bloem, F. M. J. Cozijn, W. Ubachs, and A. G. Császár, Nat. Chem. to be submitted (2023).

6F. M. J. Cozijn, P. Dupré, E. J. Salumbides, K. S. E. Eikema, and W. Ubachs, Phys. Rev. Lett., 120, 153002 (2018) DOI.

7F. Cozijn, M. Diouf, and W. Ubachs. arXiv preprint, arXiv:2303.17818 (2023), Link.

p-number: p157  Submitted on Wed, 14 Jun 2023 22:00:27 +0200